# DamageCAT: A Deep Learning Transformer Framework for Typology-Based Post-Disaster Building Damage Categorization

### v1

@@ Abstract
Natural disasters increasingly threaten communities worldwide, creating an urgent need for rapid, reliable building damage assessment to guide emergency response and recovery efforts. Current methods typically classify damage in binary (damaged/undamaged) or ordinal severity terms, limiting their practical utility. In fact, the determination of damage typology is crucial for response and recovery efforts. To address this important gap, this paper introduces DamageCAT, a novel framework that provides typology-based categorical damage descriptions rather than simple severity ratings. Accordingly, this study presents two key contributions: (1) the BD-TypoSAT dataset containing satellite image triplets (pre-disaster, post-disaster, and damage masks) from Hurricane Ida with four damage categories (partial roof damage, total roof damage, partial structural collapse, and total structural collapse), and (2) a hierarchical U-Net-based transformer architecture that effectively processes pre-post disaster image pairs to identify and categorize building damage. Despite significant class imbalances in the training data, our model achieved robust performance with overall metrics of 0.7921 Intersection over Union (IoU) and 0.8835 F1 scores across all categories. The model’s capability to recognize intricate damage typology in less common categories is especially remarkable. The DamageCAT framework advances automated damage assessment by providing actionable, typological information that better supports disaster response decision-making and resource allocation compared to traditional severity-based approaches.
@@ 1.1 Motivation & problem statement
The increasing frequency and severity of natural disasters pose significant challenges to emergency response systems worldwide. According to Gholami et al. [1], natural disasters affect approximately 350 million people annually, causing billions of dollars in damages. Climate change is expected to further exacerbate these trends, leading to more frequent and intense disasters [2]. In this context, rapid and accurate building damage assessment becomes crucial for effective disaster response and recovery. Building damage assessment serves multiple critical purposes in disaster response. As Adriano et al. [3] emphasizes, timely damage assessment is essential for both emergency response and prompt recovery efforts. Similarly, Dias et al. [4] highlights that building damage assessment is vital for guiding disaster response missions and estimating recovery needs across impacted areas. The information gathered through damage assessment enables humanitarian organizations to be prompt, effective, and efficient in supporting affected populations [5]. Traditional damage assessment methods rely heavily on manual visual interpretation of post-disaster imagery. However, this approach presents significant limitations. Alisjahbana et al. [6] point out that manual interpretation is not only time-consuming but often results in very low accuracy. The labor-intensive and tedious nature of manual annotation further hampers rapid response efforts [5]. In addition, field-based assessments may be dangerous or impossible in severely affected areas, creating information gaps in critical locations. Satellite imagery offers a promising alternative for building damage assessment. In recent years, the use of remote sensing data and machine learning has significantly enhanced damage assessment in disasters. Kahl and Chen [7] note that satellites provide a unique, comprehensive viewpoint for information retrieval from disaster-affected zones. Commercial satellite providers now enable near-daily monitoring of Earth’s surface, facilitating timely post-disaster assessments [4]. Moreover, remote sensing technologies provide rich and reliable information to support expert decision-making [5]. Despite these advantages, satellite-based damage assessment faces several challenges. Chen et al. [8] observed that the task is complicated by diverse building structures and complex environments. Liu et al. [9] highlight that building damage classification is challenging due to limited publicly available imagery data of damaged buildings and the difficulty in judging damage levels from images. Furthermore, most existing deep learning models are designed for non-damaged building segmentation and may not be suitable for damaged building classification [9]. The integration of deep learning with satellite imagery has shown promising results in automating building damage assessment. Neto and Dantas [10] demonstrated the potential of combining mathematical morphology and convolutional neural networks for building damage segmentation. Similarly, May et al. [11] utilized advanced networks such as EfficientNet and Siamese models to detect buildings and evaluate damage levels using pre- and post-event images. However, several limitations persist in current deep-learning approaches for remote sensing-based disaster-damaged assessment. Melamed et al. [12] identified bias in commonly used datasets, noting that deep learning models struggle to accurately identify isolated damaged buildings, potentially causing oversights in critical disaster scenarios and delaying humanitarian aid. In addition, Wang et al. [13] point out challenges related to extremely imbalanced datasets, where damaged buildings are significantly underrepresented compared to undamaged ones. Another significant limitation of current deep learning-based models for building damage assessment is their reliance on subjective severity scales (e.g., “low,” “medium,” “high”), which offer limited insight into the specific nature and type of damage. These severity labels often arise from coarse-grained assessments that can vary between datasets or annotators, leading to inconsistencies and ambiguity. Moreover, a single severity score cannot distinguish whether damage is confined to a roof or extends into load-bearing structures, nor can it differentiate between partial and total collapse. As a result, key nuances—such as the potential for secondary hazards or the requisite expertise for repairs—are frequently overlooked. This lack of descriptive granularity ultimately hampers targeted resource allocation, informed decision-making, and effective prioritization of rescue and recovery efforts in post-disaster scenarios.
@@
Despite the growing body of literature leveraging deep learning for post-disaster damage assessments, current severity-based classification approaches often provide only coarse estimates (e.g., “low,” “moderate,” or “high” damage). Such subjective rating scales neglect the underlying mechanisms or precise nature of the damage, limiting the applicability of these models for decision-making in real disaster scenarios. Knowing whether a structure exhibits partial roof damage or total structural collapse is not just a matter of granularity; it directly influences emergency response strategies, resource allocation, and subsequent recovery efforts. For instance, responding to a partially damaged roof may call for specialized roofing repair teams and temporary protective measures, whereas a structural collapse demands immediate search-and-rescue operations and comprehensive rebuilding plans. This study addresses that critical information gap by focusing on descriptive, type-based classifications that capture the distinct modes of structural failure, thereby enhancing both the accuracy and practical utility of automated damage assessment systems. This research context emphasizes the demand for improved approaches that reliably evaluate building damage typology across various disaster conditions, construction types, and geographic regions.
@@ 1.2 Research objectives
This study aims to advance building damage assessment by moving beyond traditional severity labels to a typology-based classification framework that captures distinct modes of structural failure. Specifically, we present a deep learning architecture that processes pre- and post-disaster satellite imagery to identify and categorize building damage with greater detail. To achieve this, we introduce BD-TypoSAT (Building Damage Typology Satellite Dataset), a novel resource containing four specific damage categories—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—offering richer representation than standard severity-based labels. We then deploy a transformer-driven, hierarchical U-Net that integrates semantic segmentation and change detection to precisely segment building footprints and classify them by damage type. This approach yields actionable information for emergency managers and stakeholders. Finally, we rigorously evaluate the model’s performance using multiple metrics (e.g., IoU, F1) to demonstrate its utility under real-world conditions, and we discuss the operational implications for disaster response, such as more informed resource allocation and prioritization of recovery efforts. By delivering detailed damage classifications—rather than broad severity ratings—our framework aims to enhance the accuracy, utility, and timeliness of post-disaster building assessments, ultimately strengthening emergency response and community resilience.
@@ 1.3 Contributions
This study makes several key contributions that collectively enhance the field of automated building damage assessment. First, we introduce the BD-TypoSAT dataset, a human-annotated resource derived from high-resolution satellite imagery captured after Hurricane Ida (2021), which departs from conventional repositories by defining four distinct damage types—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—instead of relying on binary or coarse severity labels. Second, we present a novel hierarchical U-Net-based transformer model designed to process pre- and post-disaster images and directly output typological damage classifications, offering granular and actionable insights rather than generic high/medium/low ratings. Third, we tackle the class imbalance endemic to real-world disaster data by implementing weighted loss functions (combining cross-entropy and Dice loss) and data augmentation strategies, achieving robust classification performance with an overall IoU of 0.7921 and F1 of 0.8835. In doing so, the framework provides targeted decision support, distinguishing between roof-focused damage and deeper structural failures, which ensures more effective resource allocation than severity-based methods that can often be ambiguous. Furthermore, the empirical validation on Hurricane Ida data underscores the model’s feasibility and scalability, even in complex urban environments. Lastly, by releasing both the BD-TypoSAT dataset and our transformer-driven architecture, this research lays a foundational platform for future studies aiming to expand typological damage classification to other hazards or incorporate additional data modalities, ultimately revolutionizing how post-disaster assessments and recovery efforts are conducted.
Also, this study offers novel computational contributions by integrating a hierarchical U-Net-based transformer architecture with a multi-scale difference-block mechanism that jointly processes pre- and post-disaster satellite images for fine-grained damage detection. Unlike traditional CNN pipelines, the transformer backbone facilitates long-range contextual learning, while the U-Net structure preserves critical spatial detail across multiple resolution levels. At each stage, difference blocks explicitly compare corresponding features from the two temporal images, enabling the model to pinpoint even subtle structural changes. To combat the severe class imbalance inherent in real disaster data, a weighted combination of cross-entropy and Dice losses ensures that rarer damage categories—such as total structural collapse—are not overlooked. This synergy of hierarchical feature extraction, transformer-based global context, and tailored loss weighting delivers robust, typology-specific classification performance, pushing beyond binary or severity-only assessments toward a more detailed, actionable framework for post-disaster damage analysis.
@@ 2 Related work
Research on automated building damage assessment has rapidly evolved, leveraging a wide variety of deep learning, remote sensing, and GIS-based techniques. Researchers use change detection and temporal fusion to capture pre- and post-disaster variations. For example, Zheng et al. [14] presented a deep object-based semantic change detection framework that accelerates disaster response by directly comparing pre- and post-event images. Ismail and Awad [15] proposed BLDNet, a semi-supervised framework that uses graph convolutional networks and urban domain knowledge to reduce labeling costs, even if its reliance on graph-based relationships may falter in areas with atypical urban layouts. Weber and Kané [16] fused multi-temporal satellite imagery to enhance damage assessment in dynamic environments, even though inconsistent image quality sometimes degrades its accuracy. Zhan et al. [17] employed a Siamese convolutional network to match pairwise images, boosting discriminative power while suffering when spatial alignment falters. Chen et al. [18] introduced ChangeMamba, which formulates change detection via a spatiotemporal state space model that manages uncertainties in temporal sequences, yet tuning its parameters on large datasets remains challenging. Ahn et al. [19] combined a vision foundation model with change detection to improve generalizability across disasters, although it struggles with highly specialized local scenarios. Mohammadian and Ghaderi [20] proposed Siamixformer, a fully transformer-based Siamese network that fuses temporal features for fine-grained extraction, while its high memory consumption limits its real-time application.
@@
Many studies enhance network architectures to better address damage assessment tasks. Gu et al. [21] surveyed rapid building damage identification methods based on remote sensing, emphasizing the efficiency of CNNs and transformer-based methods while noting that a lack of standardized datasets leads to inconsistent performance comparisons. Oludare et al. [22] proposed an attention-based two-stream high-resolution network tailored for building damage assessment from satellite images. Their method fuses high-resolution spatial features with attention modules to achieve strong damage identification accuracy, yet its heavy reliance on large labeled datasets limits adaptability in regions with sparse annotations. Shen et al. [23] introduced BDANet—a multi-scale CNN with cross-directional attention that effectively captures both local and global damage cues, though its high computational overhead makes it less suited for resource-constrained settings. Roohi et al. [24] combined multiple classifiers in an ensemble machine learning approach to improve flood damage predictions, even though the method requires careful tuning of numerous hyperparameters. Chen et al. [25] built a dual-task Siamese transformer framework that tackles segmentation and damage classification simultaneously; this joint optimization yields richer features but increases computational demands. Kaur et al. [26] designed a hierarchical transformer that organizes features at local, regional, and global scales to improve large-scale assessment, yet balancing depth and inference speed remains difficult. Gupta and Shah [27] introduced RescueNet, which unifies building segmentation and damage assessment into a single network to streamline training; however, the model sometimes confuses segmentation with damage levels. Some research [28, 29, 30], contributed foundational architectural modules—such as Squeeze-and-Excitation blocks, U-Net, and UNet++—that many damage assessment methods incorporate. Although these innovations boost feature recalibration and segmentation performance, they add extra parameters and memory requirements that complicate deployment in resource-limited settings.
@@
Researchers also innovate in training approaches and dataset creation to overcome the shortage of labeled data. Miglani and Kumar [31] reviewed deep learning models for hazard-damaged building detection and synthesize various strategies, noting the difficulty in transferring models across disaster types. Singh and Hoskere [32] developed a semi-supervised transformer method that leverages ultra-high-resolution aerial imagery and unlabeled data to enhance damage detection, yet it faces heavy computational loads when processing large-scale images. Ro and Gong [33] proposed a scalable method to create annotated disaster image databases, which supports subsequent model development but may underrepresent rare disaster scenarios. Xia et al. [34] pioneered a self-supervised contrastive learning approach that reduces reliance on ground-truth labels and achieves performance on par with supervised methods, even though it struggles with fine-grained damage distinctions. Finally, Gupta et al. [35] released the xBD dataset—one of the largest repositories with over 850,000 building annotations—to serve as a benchmark for damage assessment models, despite its inherent class imbalance and image quality variability.
@@
Some studies integrate diverse data sources to provide a more comprehensive picture of disaster impacts. Braik and Koliou [36] proposed a framework for automated damage assessment that integrates satellite imagery, GIS layers, and deep learning, highlighting the possibility of large-scale mapping. Their key contribution lies in combining multi-source data to improve building footprint extraction and damage classification, although the approach faces challenges in real-time deployment due to the complexity of multi-source data integration. Dong et al. [37] analyzed integrated infrastructure plans to enhance post-hazard access to critical facilities, bridging urban planning data with damage models, though it depends on extensive data availability. Techapinyawat et al. [38] applied deep learning with post-classification correction for urban land cover analysis, which can refine damage mapping but sometimes misses atypical disaster-induced changes. Gholami et al. [1] focus on the practical deployment of building damage assessment tools using satellite imagery, actively addressing issues like limited computing resources and data sharing constraints. Xue et al. [39] introduced a multi-modal approach that fuses street-view imagery with structured building data to produce refined damage predictions, even though its effectiveness is confined to regions with both reliable ground-level visuals and comprehensive metadata.
@@
Despite substantial progress in satellite-based building damage assessment, most existing methods still rely on broad severity scales—such as binary “damaged/undamaged” labels or imprecise “low/medium/high” ratings—that fail to capture the complexity of real-world damage scenarios. By focusing exclusively on severity, these approaches overlook critical descriptive details, such as whether the damage is localized to the roof or extends into load-bearing structures—information that can be pivotal for determining repair priority, resource allocation, and insurance decisions. Moreover, nuanced indicators like partial collapses or roof breaches can signal specific failure mechanisms and potential secondary hazards, yet these remain unaddressed in standard severity-centric frameworks. As a result, decision-makers—ranging from emergency responders to policy planners—lack the granular insights needed to tailor interventions and expedite recovery. Addressing this shortfall necessitates integrated frameworks that move beyond measuring “how much” damage occurred to clarifying “what kind” of damage has taken place, enabling more targeted, context-aware strategies for disaster response, reconstruction, and long-term resilience planning.

### v2

@@ Abstract
Rapid, accurate, and descriptive building damage assessment is critical for directing post-disaster resources, yet current automated methods typically provide only binary (damaged/undamaged) or ordinal severity scales. This paper introduces DamageCAT, a framework that advances damage assessment through typology-based categorical classifications. We contribute: (1) the BD-TypoSAT dataset containing satellite image triplets from Hurricane Ida with four damage categories—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—and (2) a hierarchical U-Net-based transformer architecture for processing pre- and post-disaster image pairs. Our model achieves 0.737 IoU and 0.846 F1-score overall, with cross-event evaluation demonstrating transferability across Hurricane Harvey, Florence, and Michael data. While performance varies across damage categories due to class imbalance, the framework shows that typology-based classification can provide more actionable damage assessments than traditional severity-based approaches, enabling targeted emergency response and resource allocation.
@@ 1.1 Motivation & problem statement
The increasing frequency and severity of natural disasters pose significant challenges to emergency response systems worldwide. According to Gholami et al. [1], natural disasters affect approximately 350 million people annually, causing billions of dollars in damages. Climate change is expected to further exacerbate these trends, leading to more frequent and intense disasters [2]. In this context, rapid and accurate building damage assessment becomes crucial for effective disaster response and recovery. Building damage assessment serves multiple critical purposes in disaster response. As Adriano et al. [3] emphasizes, timely damage assessment is essential for both emergency response and prompt recovery efforts. Similarly, Dias et al. [4] highlights that building damage assessment is vital for guiding disaster response missions and estimating recovery needs across impacted areas. The information gathered through damage assessment enables humanitarian organizations to be prompt, effective, and efficient in supporting affected populations [5]. Traditional damage assessment methods rely heavily on manual visual interpretation of post-disaster imagery. However, this approach presents significant limitations. Alisjahbana et al. [6] point out that manual interpretation is not only time-consuming but often results in very low accuracy. The labor-intensive and tedious nature of manual annotation further hampers rapid response efforts [5]. In addition, field-based assessments may be dangerous or impossible in severely affected areas, creating information gaps in critical locations. Satellite imagery offers a promising alternative for building damage assessment. In recent years, the use of remote sensing data and machine learning has significantly enhanced damage assessment in disasters. Kahl and Chen [7] note that satellites provide a unique, comprehensive viewpoint for information retrieval from disaster-affected zones. Commercial satellite providers now enable near-daily monitoring of Earth’s surface, facilitating timely post-disaster assessments [4]. Moreover, remote sensing technologies provide rich and reliable information to support expert decision-making [5]. Despite these advantages, satellite-based damage assessment faces several challenges. Chen et al. [8] observed that the task is complicated by diverse building structures and complex environments. Liu et al. [9] highlight that building damage classification is challenging due to limited publicly available imagery data of damaged buildings and the difficulty in judging damage levels from images. Furthermore, most existing deep learning models are designed for non-damaged building segmentation and may not be suitable for damaged building classification [9]. The integration of deep learning with satellite imagery has shown promising results in automating building damage assessment. Neto and Dantas [10] demonstrated the potential of combining mathematical morphology and convolutional neural networks for building damage segmentation. Similarly, May et al. [11] utilized advanced networks such as EfficientNet and Siamese models to detect buildings and evaluate damage levels using pre- and post-event images.
However, several limitations persist in current deep-learning approaches for remote sensing-based disaster-damaged assessment. Melamed et al. [12] identified bias in commonly used datasets, noting that deep learning models struggle to accurately identify isolated damaged buildings, potentially causing oversights in critical disaster scenarios and delaying humanitarian aid. In addition, Wang et al. [13] point out challenges related to extremely imbalanced datasets, where damaged buildings are significantly underrepresented compared to undamaged ones. Another significant limitation of current deep learning-based models for building damage assessment is their reliance on subjective severity scales (e.g., “low,” “moderate,” “high”), which offer limited insight into the specific nature and type of damage. These severity labels often arise from coarse-grained assessments that can vary between datasets or annotators, leading to inconsistencies and ambiguity. Moreover, a single severity score cannot distinguish whether damage is confined to a roof or extends into load-bearing structures, nor can it differentiate between partial and total collapse. As a result, key nuances—such as the potential for secondary hazards or the requisite expertise for repairs—are frequently overlooked. This lack of descriptive granularity ultimately hampers targeted resource allocation, informed decision-making, and effective prioritization of rescue and recovery efforts in post-disaster scenarios.
@@
Despite the growing amount of literature leveraging deep learning for post-disaster damage assessments, current severity-based classification approaches often provide only coarse estimates. While previous studies have made important strides in damage classification, subjective rating scales can neglect the underlying mechanisms or precise nature of the damage, potentially limiting model applicability for decision-making in real disaster scenarios. Knowing whether a structure exhibits partial roof damage or total structural collapse is not just a matter of granularity; it directly influences emergency response strategies, resource allocation, and subsequent recovery efforts. For instance, responding to a partially damaged roof may call for specialized roofing repair teams and temporary protective measures, whereas a structural collapse demands immediate search-and-rescue operations and comprehensive rebuilding plans. Building upon foundational work in remote sensing-based damage assessment, this study explores a typology-based classification approach that aims to capture distinct modes of structural failure, complementing existing automated assessment systems. This research context emphasizes opportunities to enhance damage evaluation across various disaster conditions, construction types, and geographic regions, while acknowledging the inherent challenges of satellite-based assessment.
@@ 1.2 Research objectives
This study aims to contribute to building damage assessment by exploring typology-based classification framework that captures distinct modes of structural failure. We present a deep learning architecture that processes pre- and post-disaster satellite imagery to identify and categorize building damage with greater detail. To support this exploration, we introduce BD-TypoSAT (Building Damage Typology Satellite Dataset), containing four specific damage categories—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—offering potentially richer representation than standard severity-based labels. We then deploy a transformer-driven, hierarchical U-Net that integrates semantic segmentation and change detection to segment building footprints and classify them by damage type. This approach aims to yield actionable information for emergency managers and stakeholders. Finally, we rigorously evaluate the model’s performance using multiple metrics (e.g., IoU, F1) to demonstrate its utility under real-world conditions, and we discuss both the operational implications and limitations for disaster response, such as more informed resource allocation and prioritization of recovery efforts.
@@ 1.3 Contributions
This study makes several key contributions to automated building damage assessment. We introduce the BD-TypoSAT dataset, derived from high-resolution satellite imagery captured after Hurricane Ida (2021), which categorizes building damage into four distinct typological classes—partial roof damage, total roof damage, partial structural collapse, and total structural collapse. This typological approach offers a more nuanced alternative to traditional binary or coarse severity classifications, enabling targeted decision support by distinguishing between surface-level roof damage and deeper structural failures.
Our computational framework integrates a hierarchical U-Net-based transformer architecture with multi-scale difference blocks that jointly process pre- and post-disaster satellite images. The transformer backbone facilitates long-range contextual learning while preserving critical spatial detail through the U-Net structure. At each resolution level, difference blocks explicitly compare corresponding features between temporal images, enabling detection of subtle structural changes. To address the severe class imbalance inherent in disaster data, we implement a weighted combination of cross-entropy and Dice losses, ensuring that rarer damage categories such as total structural collapse are not overlooked during training.
Quantitative validation demonstrates promising performance with an overall IoU of 0.737 and F1-score of 0.846, while cross-event evaluation on Hurricanes Harvey, Florence, and Michael shows encouraging transferability across different disaster contexts. Though accuracy varies across urban environments and damage types, the framework provides actionable insights for emergency response applications. By releasing both the BD-TypoSAT dataset and our transformer-driven architecture, this research contributes valuable resources for future studies aiming to expand typological damage classification to other hazards or incorporate additional data modalities.
@@ 2 Related work
Research on automated building damage assessment has rapidly evolved, leveraging a wide variety of remote sensing, GIS-based techniques, and deep learning.
In the pre-deep learning era, remote sensing imagery was recognized as a valuable tool for rapid post-disaster damage mapping, allowing for swift assessment over wide areas [14], offering advantages like efficiency and cost-effectiveness compared to field surveys. For detailed structural damage assessment at the individual building level, Womble et al. [15] highlighted the necessity of very high spatial resolution imagery, typically sub-meter pixel size, enabling the generation of various map products focusing on individual buildings, city blocks, or aggregated areas, often through visual interpretation by experienced analysts. However, this foundational approach faced significant limitations. Kerle [14] noted that imagery quality can be poor or insufficient, and clouds frequently obscure affected areas. The predominantly vertical perspective of satellite imagery, as pointed out by Yamazaki et al. [16], makes damage visible primarily on roofs, often missing facade damage or interior issues, especially less severe ones, and making it difficult to distinguish between different damage grades. This ambiguity was further compounded by a lack of universally accepted nomenclature for image-mapped damage [14]. Moreover, accurately assessing map quality proved challenging due to a scarcity of suitable ground truth data; while field surveys offer higher accuracy, they can be incomplete or inconsistent, particularly when conducted rapidly [14], and obtaining ground truth at a scale comparable to high-resolution imagery interpretation remains difficult [17]. These early challenges emphasized the need for more robust methods and data sources.
To address some of these limitations, integrating different data sources was explored to improve damage mapping. The use of GIS became crucial for integrating remote sensing data with other information, such as pre-disaster building databases [18, 15], and combining image-based mapping with limited ground validation was considered an ideal approach, particularly for areas difficult to interpret from imagery alone. Kerle [14] emphasized that using multi-temporal imagery (pre- and post-event) is often more effective for change detection related to damage than relying on post-event images alone, leading to the exploration of automated change detection techniques like threshold segmentation and change vector analysis. The development of multimodal datasets, for instance combining optical and SAR imagery, also began to support training more advanced algorithms [14]. However, even with these advancements, challenges related to data perspective persisted. As noted by Womble et al. [15] and Yamazaki et al. [16], damage assessment primarily relied on nadir (vertical) perspectives, making facade damage difficult to discern. While airborne platforms offered potential for oblique views, a detailed review of analysis techniques for oblique satellite or aerial imagery remained limited in early literature [15], representing a gap in comprehensive damage visibility.
@@
With the advent of deep learning, researchers increasingly utilized change detection and temporal fusion techniques to capture pre- and post-disaster variations. For instance, Zheng et al. [19] presented a deep object-based semantic change detection framework to accelerate disaster response by directly comparing pre- and post-event images. To address labeling costs, Ismail and Awad [20] proposed BLDNet, a semi-supervised framework using graph convolutional networks and urban domain knowledge, though its reliance on graph-based relationships may falter in areas with atypical urban layouts. Weber and Kané [21] focused on fusing multi-temporal satellite imagery to enhance damage assessment in dynamic environments, but inconsistent image quality could sometimes degrade accuracy. Earlier, Zhan et al. [22] employed a Siamese convolutional network to match pairwise images, boosting discriminative power, yet this approach could suffer when spatial alignment faltered. More recently, Chen et al. [23] introduced ChangeMamba, formulating change detection via a spatiotemporal state space model to manage uncertainties in temporal sequences; however, tuning its parameters on large datasets remains challenging. In an effort to improve generalizability, Ahn et al. [24] combined a vision foundation model with change detection, although it was found to struggle with highly specialized local scenarios. Mohammadian and Ghaderi [25] proposed Siamixformer, a fully transformer-based Siamese network that fuses temporal features for fine-grained extraction, but its high memory consumption can limit real-time application. These approaches demonstrate significant progress, yet often come with trade-offs in complexity, generalizability, or computational cost.
@@
Beyond temporal fusion, many studies have focused on enhancing network architectures to better address damage assessment tasks. Gu et al. [26] surveyed rapid building damage identification methods, emphasizing the efficiency of CNNs and transformer-based approaches, while critically noting that a lack of standardized datasets leads to inconsistent performance comparisons. Specific architectural innovations include the work of Oludare et al. [27], who proposed an attention-based two-stream high-resolution network for satellite images; their method fuses high-resolution spatial features with attention modules for strong accuracy, but its heavy reliance on large labeled datasets limits adaptability in sparsely annotated regions. Shen et al. [28] introduced BDANet, a multi-scale CNN with cross-directional attention to capture local and global damage cues, though its high computational overhead poses challenges for resource-constrained settings. An ensemble machine learning approach by Roohi et al. [29] combined multiple classifiers to improve flood damage predictions, yet this method requires careful tuning of numerous hyperparameters. Chen et al. [30] developed a dual-task Siamese transformer framework for simultaneous segmentation and damage classification, which, while yielding richer features through joint optimization, also increases computational demands. Similarly, Kaur et al. [31] designed a hierarchical transformer for large-scale assessment by organizing features at multiple scales, but balancing depth and inference speed remains difficult. Gupta and Shah [32] introduced RescueNet to unify building segmentation and damage assessment, streamlining training, though the model sometimes confused segmentation with damage levels. Foundational architectural modules such as Squeeze-and-Excitation blocks [33], U-Net [34], and UNet++ [35] have been widely incorporated, boosting feature recalibration and segmentation; however, these innovations often add parameters and memory requirements, complicating deployment in resource-limited settings.
@@
Recognizing the critical challenge of labeled data scarcity, researchers have also innovated in training approaches and dataset creation. Miglani and Kumar [36] reviewed deep learning models for hazard-damaged building detection, synthesizing various strategies and highlighting the difficulty in transferring models across different disaster types. To mitigate data limitations, Singh and Hoskere [37] developed a semi-supervised transformer method leveraging ultra-high-resolution aerial imagery and unlabeled data, though it faces heavy computational loads with large-scale images. Ro and Gong [38] proposed a scalable method for creating annotated disaster image databases, which, while supporting model development, may underrepresent rare disaster scenarios. A self-supervised contrastive learning approach pioneered by Xia et al. [39] reduces reliance on ground-truth labels, achieving performance comparable to supervised methods, but it can struggle with fine-grained damage distinctions. The release of datasets like xBD by Gupta et al. [40], one of the largest repositories with over 850,000 building annotations, has been crucial for benchmarking models, despite its inherent class imbalance and image quality variability. These efforts aim to make deep learning more applicable in data-scarce disaster contexts, though challenges in model transferability and dataset representativeness persist.
@@
Further efforts to enhance damage assessment involve integrating diverse data sources for a more comprehensive picture of disaster impacts. For example, Braik and Koliou [41] proposed a framework integrating satellite imagery, GIS layers, and deep learning for large-scale mapping, improving building footprint extraction and damage classification; however, this approach faces challenges in real-time deployment due to multi-source data integration complexity. Dong et al. [42] analyzed integrated infrastructure plans to enhance post-hazard access to critical facilities, bridging urban planning data with damage models, though this method depends on extensive data availability. Techapinyawat et al. [43] applied deep learning with post-classification correction for urban land cover analysis, which can refine damage mapping but may sometimes miss atypical disaster-induced changes. Focusing on practical deployment, Gholami et al. [1] addressed issues like limited computing resources and data sharing constraints for satellite imagery-based tools. Xue et al. [44] introduced a multi-modal approach fusing street-view imagery with structured building data for refined damage predictions, though its effectiveness is confined to regions with both reliable ground-level visuals and comprehensive metadata. While these integrated approaches offer richer insights, they often introduce complexities in data fusion, real-time processing, and data availability.
@@
Despite the substantial progress reviewed, a significant gap persists in satellite-based building damage assessment: most existing methods still rely on broad severity scales, such as binary “damaged/undamaged” labels or imprecise “low/medium/high” ratings. As highlighted throughout the literature, these approaches often fail to capture the complexity of real-world damage scenarios. By focusing predominantly on severity, they can overlook critical descriptive details, such as whether damage is localized to the roof or extends into load-bearing structures—information pivotal for repair priority, resource allocation, and insurance decisions. Nuanced indicators like partial collapses or roof breaches, which can signal specific failure mechanisms and potential secondary hazards, often remain unaddressed in standard severity-centric frameworks. Consequently, decision-makers, from emergency responders to policy planners, may lack the granular insights needed to tailor interventions and expedite recovery. Addressing this shortfall necessitates a shift towards frameworks that move beyond merely measuring “how much” damage occurred to clarifying “what kind” of damage has taken place, enabling more targeted, context-aware strategies for disaster response, reconstruction, and long-term resilience planning. This study aims to contribute to filling this gap by exploring a typology-based classification approach.